# On the equivalence of NMDS codes

### v1

@@ 1 Introduction
By adding suitable projective points to the $n$-arcs in $\mathrm{PG}(k-1,q)$, many NMDS codes of dimension $k$ have been constructed. For the case $k=3$, see [40, 20, 21, 42, 13]. For the cases $k=4$ and $k=5$, see [41, 15, 11]. Moreover, there are many NMDS codes with the same parameters derived from these constructions. In this paper, we investigate the monomial equivalence problem for such NMDS codes and provide some new constructions. Specifically, by utilizing hyperovals in $\mathrm{PG}(2,q)$, we propose an infinite family of NMDS codes with length $q+3$ and dimension $3$ for even $q$ (see Theorem 3.9). Through detailed analysis of the action of the homography stabilizers of distinct hyperovals on the points not in hyperovals, we fully determine the equivalence classes of these NMDS codes. By analyzing the orbits of points and planes of $\mathrm{PG}(3,q)$ under the stabilizers of arcs, we provide a unified geometric perspective for constructing several equivalence classes of NMDS codes with length $q+2$ and dimension $4$, derived from arcs in $\mathrm{PG}(3,q)$ with additional points (see Theorems 4.26 and 4.30 ).
@@ 2.2 Hyperovals
An $n$-arc in $\mathrm{PG}(2,q)$ is a set of $n$ points such that no three are collinear. It is know that $n\leq q+2$, with equality holding if and only if $q$ is even. Let $q=2^{m}$ for a positive integer $m$. A hyperoval $\mathcal{HO}$ in $\mathrm{PG}(2,q)$ is an arc with $q+2$ points, where every line in $\mathrm{PG}(2,q)$ intersects $\mathcal{HO}$ in either 0 or $2$ points. Two hyperovals are equivalent if one can be mapped to the other via an automorphism of $\mathrm{PG}(2,q)$. All hyperovals in $\mathrm{PG}(2,q)$ can be constructed using a special type of permutation polynomials over $\mathbb{F}_{q}$, as described in the following Theorem [16].
@@
By [6], for even $q$, a conic together with its nucleus (the intersection point of its tangents) forms a hyperoval. Such a hyperoval is said to be regular. As shown by [34], for $q=2,4,8$, every hyperoval is regular. However, for $q=2^{m}$ with $m\geq 4$, irregular hyperovals exist. Any polynomial satisfying the two conditions in Theorem 2.3 is called an oval polynomial (or o-polynomial for short). Table 1 gives a list of all known o-polynomials over $\mathbb{F}_{q}$ with $q=2^{m}$, also see [18, Table 3]. Here,
@@
[28, Lemma 4.6.6] If hyperovals $K$ and $K^{\prime}$ in $\mathrm{PG}(2,q)$ have $\frac{1}{2}(q+2)+n$ points in common, with $n>0$, then $K=K^{\prime}$.
@@
The proof is similar to that of [42, Lemma 3.1], and so we omit the details. ∎
@@ 2.3 Arcs in $\mathrm{PG}(3,q)$
An arc in $\mathrm{PG}(3,q)$, $q\geq 4$ has at most $q+1$ points [17, Theorems 21.2.4 and 21.3.8]. If $q=2^{m}$, a $(q+1)$-arc in $\mathrm{PG}(3,q)$ is projectively equivalent to the point-set
@@
where $\gcd(m,h)=1$ [17, Theorem 21.3.15]. For $h=1$, $\mathcal{S}_{1}$ is called a twisted cubic. When $q$ is odd, every $(q+1)$-arc in $\mathrm{PG}(3,q)$ is a twisted cubic. A line in $\mathrm{PG}(3,q)$ joining two distinct points of $\mathcal{S}_{h}$ is called a real chord. Note that no four points of $\mathcal{S}_{h}$ are coplanar. Therefore, each real chord contains exactly two points of $\mathcal{S}_{h}$. Let $R$ denote the set of points lying on real chords but not in $\mathcal{S}_{h}$. Since there are $q(q+1)/2$ real chords, it follows that $|R|=(q^{3}-q)/2$. The planes $\pi(P_{x})$ and $\pi(P_{\infty})$, defined by the equations $x^{2^{h}+1}X_{1}+x^{2^{h}}X_{2}+xX_{3}+X_{4}=0$ and $X_{1}=0$ are called the osculating planes of $\mathcal{S}_{h}$ at $P_{x}$ and $P_{\infty}$, respectively. The osculating plane $\pi(P)$ intersects $\mathcal{S}_{h}$ only at point $P$. We end this section with the following lemma.
@@
[37, Theorem A] Let $K$ be a $k$-arc of $\mathrm{PG}(3,q)$, $q$ even and $q\neq 2$. Assume that $k>q-\sqrt{q}/2+9/4$. Then $K$ lies on a unique $(q+1)$-arc.
@@ 3.1 NMDS codes from hyperovals
where $\mathbf{g}_{q+3}=(a,b,c)^{\mathrm{T}}\neq\mathbf{0}$, and $\mathbf{g}_{q+3}\neq\lambda\mathbf{g}_{i}$ for some $\lambda\in\mathbb{F}^{*}_{q}$ and any $i\in\{1,\dots,q+2\}$. For convenience, we identify each vector $\mathbf{g}_{i}$ with the point in $\mathrm{PG}(2,q)$ spanned by $\mathbf{g}_{i}$. This ensures that the points $\mathbf{g}_{i}$ are distinct for $i=1,\dots,q+3$. Let $\mathcal{C}$ be the linear code over $\mathbb{F}_{q}$ with generator matrix $G$. For any $\mathbf{g}_{q+3}$ , we first prove the following theorem, which generalizes [40, Theorem 8] and [21, Theorem 10].
@@ 3.2 Monomial equivalence classes of NMDS codes
[16, Corollary 6][28, Theorem 4.11.4] A regular hyperoval in $\mathrm{PG}(2,q)$, $q>4$ even, has homography stabilizer of order $q(q^{2}-1)$ isomorphic to $\mathrm{PGL}(2,q)$. This stabilizer consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
@@
For regular o-polynomial, by setting $\mathbf{g}_{q+3}$ as $(0,1,1)^{\mathrm{T}}$ and $(1,1,0)^{\mathrm{T}}$ in (3.1), we obtain two distinct matrices $G_{1}$ and $G_{2}$. Two $[q+3,3,q]$ NMDS codes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ with generator matrices $G_{1}$ and $G_{2}$, were constructed in [40, Theorem 8] and [21, Theorem 10], respectively. By Theorem 3.13, we see that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are monomially equivalent.
Two [q+1,3,q2][q+1,3,q-2] NMDS codes $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ with the same weight enumerator were provided in [40, Theorem 10] and [21, Theorem 13] respectively, where $q=2^{m}$ and $m\geq 3$ is odd. The generator matrices $G_{i}$ of $\mathcal{C}_{i}$, for $i=3,4$, are as follows:
@@
[25, Theorem 4.3][28, Theorem 4.11.6] An irregular translation hyperoval with o-polynomial $f(x)=x^{\alpha}$ in $\mathrm{PG}(2,q)$, where $\alpha=2^{h},q=2^{m}\geq 16$, and $\gcd(h,m)=1$, has homography stabilizer of order q(q1)q(q-1). This stabilizer is isomorphic to $\mathrm{AGL}(1,q)$ and consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
@@
[25, Theorem 4.4][28, Theorem 4.12.8] Let $\mathcal{HO}$ be a hyperoval with o-polynomial $f(x)=x^{n}$, but not a translation hyperoval. If $n^{2}-n+1\not\equiv 0\pmod{q-1}$, then $\mathcal{HO}$ has homography stabilizer $H$ of order (q1)(q-1) which consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
@@
By [24, Theroem 3.2], the O’Keefe-Penttila hyperoval has homography stabilizer of order $3$, generated by the matrix $\left(\begin{array}[]{ccc}0&1&0\\ 1&1&1\\ 0&0&1\end{array}\right)$. It cannot fix any point not in the hyperoval. The claim then follows from Corollary 3.11. ∎
@@ 4.1 $q$ even
where $a,b,c,d\in\mathbb{F}_{q}$ satisfy $ad+bc=1$, see [17, Lemma 21.3.13, Theorem 21.3.17]. We now present the results on the orbits of points and planes under the action of $K_{h}$.
@@
[8, Proposition 3.2] The group $K_{h}$ has five orbits on the points of $\mathrm{PG}(3,q)$:
@@
[8, Proposition 3.4] The group $K_{h}$ has five orbits on the planes of $\mathrm{PG}(3,q)$:
@@
By Lemma 4.24, if $q\equiv-1\pmod{3}$ and $\mathbf{g}_{q+2}\in\mathcal{M}_{4}$ or $q\equiv 1\pmod{3}$ and $\mathbf{g}_{q+2}\in\mathcal{M}_{3}\cup\mathcal{M}_{5}$, then the corresponding linear code is not an NMDS code. For the remaining cases, we only need consider the number of nonzero vectors $u\in\mathbb{F}^{4}_{q}$ such that $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=4$. Suppose that a nonzero vector $u\in\mathbb{F}^{4}_{q}$ satisfies this condition. Then $\mathbf{g}_{q+2}\in\mathcal{H}_{u}$ and so $|\mathcal{H}_{u}\cap\mathcal{S}_{h}|=3$. By Lemma 4.25, all planes containing exactly three points of $\mathcal{S}_{h}$ form a $K_{h}$-orbit $\pi(\mathcal{M}_{3})$. This means the number of nonzero vectors $u\in\mathbb{F}^{4}_{q}$ such that $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=4$ is equal to q1q-1 times the number of planes in $\pi(\mathcal{M}_{3})$ through the point $\mathbf{g}_{q+2}$. By [8, Table 1], these numbers are all great than 0 and only depend on the $K_{h}$-orbit on points to which $\mathbf{g}_{q+2}$ belongs. Thus, the remaining linear codes are all NMDS codes. Obviously, the codes corresponding to $\mathbf{g}_{q+2}$ in the same $K_{h}$-orbit are monomially equivalent, while the codes corresponding to $\mathbf{g}_{q+2}$ in different $K_{h}$-orbits are not monomially equivalent as they have different numbers of minimum weight codes. Therefore, there are two or three equivalence classes of NMDS codes in $\mathscr{C}$ depending on whether $q\equiv 1\pmod{3}$ or $q\equiv-1\pmod{3}$. Finally, by Lemma 2.2, the weight distributions of these codes can be fully determined and we provide them in Table 2. ∎
@@
Let $\mathbf{g}^{\mathrm{T}}_{q+2}=(0,0,1,0)\in\mathcal{M}_{2}$ in (4.1). Then we can directly obtain the corresponding result from [11, Theorem 15] using Theorem 4.26.
@@ 4.2 $q$ odd
[17, Lemma 21.1.3][4, Theorem 2.2] Let $q\equiv 0\pmod{3}$. The group $K_{q}$ has five orbits on the points of $\mathrm{PG}(3,q)$:
@@ 4.3 NMDS codes with parameters [q+3,4,q1][q+3,4,q-1]
where $\mathbb{F}^{*}_{q}=\mathbb{F}_{q}\setminus\{0\}=\{\alpha_{1},\dots,\alpha_{q-1}\}$ and $\mathbf{g}_{q+2},\mathbf{g}_{q+3}\neq\mathbf{0}$. By selecting specific $\mathbf{g}_{q+2}$ and $\mathbf{g}_{q+3}$, we can construct certain NMDS codes. For example, let $p=2$ and $h$ be a positive integer satisfying $\gcd(m,h)=1$, where $m\geq 3$ is odd. Take $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,0,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ or $\mathbf{g}_{q+3}=(0,1,1,0)^{\mathrm{T}}\in\mathcal{M}_{3}$, where $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$ are the point-orbits of $K_{h}$ defined in Lemma 4.24. Then the linear code generated by $G_{3}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code (see [11, Theorems 18 and 19]). A simpler proof of this result can be provided via Theorem 4.26.
@@
Let $h=1$ and $\gcd(3,q-1)=1$. Taking $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,0,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ in (4.2), the linear code generated by $G_{3}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code (see [41, Theorem 9]). Similarly to the proof of Theorem 4.31, we further have the following Theorem from Tables 2 and 3.
@@
When $q=2^{m}$ with $m$ odd, Theorems 4.31 and 4.33 show that $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$ have identical weight enumerators by Lemma 2.2. However, as noted in [11, Remark 22], $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$ are not monomially equivalent [q+3,4,q1][q+3,4,q-1] NMDS codes for $q=2^{3},2^{5},2^{7}$ and some possible values of $h$. Finally, we have the following proposition.

### v2

@@ 1 Introduction
By adding suitable projective points to the $n$-arcs in $\mathrm{PG}(k-1,q)$, many NMDS codes of dimension $k$ have been constructed. For the case $k=3$, see [41, 21, 22, 43, 14]. For the cases $k=4$ and $k=5$, see [42, 16, 12, 8]. Moreover, there are many NMDS codes with the same parameters derived from these constructions. In this paper, we investigate the monomial equivalence problem for such NMDS codes and provide some new constructions. Specifically, by utilizing hyperovals in $\mathrm{PG}(2,q)$, we propose an infinite family of NMDS codes with length $q+3$ and dimension $3$ for even $q$ (see Theorem 3.9). Through detailed analysis of the action of the homography stabilizers of distinct hyperovals on the points not in hyperovals, we fully determine the equivalence classes of these NMDS codes. By analyzing the orbits of points and planes of $\mathrm{PG}(3,q)$ under the stabilizers of arcs, we provide a unified geometric perspective for constructing several equivalence classes of NMDS codes with length $q+2$ and dimension $4$, derived from arcs in $\mathrm{PG}(3,q)$ with additional points (see Theorems 4.26 and 4.30 ).
@@ 2.2 Hyperovals
An $n$-arc in $\mathrm{PG}(2,q)$ is a set of $n$ points such that no three are collinear. It is known that $n\leq q+2$, with equality holding if and only if $q$ is even. Let $q=2^{m}$ for a positive integer $m$. A hyperoval $\mathcal{HO}$ in $\mathrm{PG}(2,q)$ is an arc with $q+2$ points, where every line in $\mathrm{PG}(2,q)$ intersects $\mathcal{HO}$ in either 0 or $2$ points. Two hyperovals are equivalent if one can be mapped to the other via an automorphism of $\mathrm{PG}(2,q)$. All hyperovals in $\mathrm{PG}(2,q)$ can be constructed using a special type of permutation polynomials over $\mathbb{F}_{q}$, as described in the following Theorem [17].
@@
By [6], for even $q$, a conic together with its nucleus (the intersection point of its tangents) forms a hyperoval. Such a hyperoval is said to be regular. As shown by [35], for $q=2,4,8$, every hyperoval is regular. However, for $q=2^{m}$ with $m\geq 4$, irregular hyperovals exist. Any polynomial satisfying the two conditions in Theorem 2.3 is called an oval polynomial (or o-polynomial for short). Table 1 gives a list of all known o-polynomials over $\mathbb{F}_{q}$ with $q=2^{m}$, also see [19, Table 3]. Here,
@@
[29, Lemma 4.6.6] If hyperovals $K$ and $K^{\prime}$ in $\mathrm{PG}(2,q)$ have $\frac{1}{2}(q+2)+n$ points in common, with $n>0$, then $K=K^{\prime}$.
@@
The proof is similar to that of [43, Lemma 3.1], and so we omit the details. ∎
@@ 2.3 Arcs in $\mathrm{PG}(3,q)$
An arc in $\mathrm{PG}(3,q)$, $q\geq 4$ has at most $q+1$ points [18, Theorems 21.2.4 and 21.3.8]. If $q=2^{m}$, a $(q+1)$-arc in $\mathrm{PG}(3,q)$ is projectively equivalent to the point-set
@@
where $\gcd(m,h)=1$ [18, Theorem 21.3.15]. For $h=1$, $\mathcal{S}_{1}$ is called a twisted cubic. When $q$ is odd, every $(q+1)$-arc in $\mathrm{PG}(3,q)$ is a twisted cubic. A line in $\mathrm{PG}(3,q)$ joining two distinct points of $\mathcal{S}_{h}$ is called a real chord. Note that no four points of $\mathcal{S}_{h}$ are coplanar. Therefore, each real chord contains exactly two points of $\mathcal{S}_{h}$. Let $R$ denote the set of points lying on real chords but not in $\mathcal{S}_{h}$. Since there are $q(q+1)/2$ real chords, it follows that $|R|=(q^{3}-q)/2$. The planes $\pi(P_{x})$ and $\pi(P_{\infty})$, defined by the equations $x^{2^{h}+1}X_{1}+x^{2^{h}}X_{2}+xX_{3}+X_{4}=0$ and $X_{1}=0$ are called the osculating planes of $\mathcal{S}_{h}$ at $P_{x}$ and $P_{\infty}$, respectively. The osculating plane $\pi(P)$ intersects $\mathcal{S}_{h}$ only at point $P$. We end this section with the following lemma.
@@
[38, Theorem A] Let $K$ be a $k$-arc of $\mathrm{PG}(3,q)$, $q$ even and $q\neq 2$. Assume that $k>q-\sqrt{q}/2+9/4$. Then $K$ lies on a unique $(q+1)$-arc.
@@ 3.1 NMDS codes from hyperovals
where $\mathbf{g}_{q+3}=(a,b,c)^{\mathrm{T}}\neq\mathbf{0}$, and $\mathbf{g}_{q+3}\neq\lambda\mathbf{g}_{i}$ for some $\lambda\in\mathbb{F}^{*}_{q}$ and any $i\in\{1,\dots,q+2\}$. For convenience, we identify each vector $\mathbf{g}_{i}$ with the point in $\mathrm{PG}(2,q)$ spanned by $\mathbf{g}_{i}$. This ensures that the points $\mathbf{g}_{i}$ are distinct for $i=1,\dots,q+3$. Let $\mathcal{C}$ be the linear code over $\mathbb{F}_{q}$ with generator matrix $G$. For any $\mathbf{g}_{q+3}$, we first prove the following theorem, which generalizes [41, Theorem 8] and [22, Theorem 10].
@@ 3.2 Monomial equivalence classes of NMDS codes
[17, Corollary 6][29, Theorem 4.11.4] A regular hyperoval in $\mathrm{PG}(2,q)$, $q>4$ even, has homography stabilizer of order $q(q^{2}-1)$ isomorphic to $\mathrm{PGL}(2,q)$. This stabilizer consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
@@
For regular o-polynomial, by setting $\mathbf{g}_{q+3}$ as $(0,1,1)^{\mathrm{T}}$ and $(1,1,0)^{\mathrm{T}}$ in (3.1), we obtain two distinct matrices $G_{1}$ and $G_{2}$. Two $[q+3,3,q]$ NMDS codes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ with generator matrices $G_{1}$ and $G_{2}$, were constructed in [41, Theorem 8] and [22, Theorem 10], respectively. By Theorem 3.13, we see that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are monomially equivalent.
Two [q+1,3,q2][q+1,3,q-2] NMDS codes $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ with the same weight enumerator were provided in [41, Theorem 10] and [22, Theorem 13] respectively, where $q=2^{m}$ and $m\geq 3$ is odd. The generator matrices $G_{i}$ of $\mathcal{C}_{i}$, for $i=3,4$, are as follows:
@@
[26, Theorem 4.3][29, Theorem 4.11.6] An irregular translation hyperoval with o-polynomial $f(x)=x^{\alpha}$ in $\mathrm{PG}(2,q)$, where $\alpha=2^{h},q=2^{m}\geq 16$, and $\gcd(h,m)=1$, has homography stabilizer of order q(q1)q(q-1). This stabilizer is isomorphic to $\mathrm{AGL}(1,q)$ and consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
@@
[26, Theorem 4.4][29, Theorem 4.12.8] Let $\mathcal{HO}$ be a hyperoval with o-polynomial $f(x)=x^{n}$, but not a translation hyperoval. If $n^{2}-n+1\not\equiv 0\pmod{q-1}$, then $\mathcal{HO}$ has homography stabilizer $H$ of order (q1)(q-1) which consists of the maps: $(x,y,z)^{\mathrm{T}}\mapsto A(x,y,z)^{\mathrm{T}}$, where
@@
By [25, Theroem 3.2], the O’Keefe-Penttila hyperoval has homography stabilizer of order $3$, generated by the matrix $\left(\begin{array}[]{ccc}0&1&0\\ 1&1&1\\ 0&0&1\end{array}\right)$. It cannot fix any point not in the hyperoval. The claim then follows from Corollary 3.11. ∎
@@ 4.1 $q$ even
where $a,b,c,d\in\mathbb{F}_{q}$ satisfy $ad+bc=1$, see [18, Lemma 21.3.13, Theorem 21.3.17]. We now present the results on the orbits of points and planes under the action of $K_{h}$.
@@
[9, Proposition 3.2] The group $K_{h}$ has five orbits on the points of $\mathrm{PG}(3,q)$:
@@
[9, Proposition 3.4] The group $K_{h}$ has five orbits on the planes of $\mathrm{PG}(3,q)$:
@@
By Lemma 4.24, if $q\equiv-1\pmod{3}$ and $\mathbf{g}_{q+2}\in\mathcal{M}_{4}$ or $q\equiv 1\pmod{3}$ and $\mathbf{g}_{q+2}\in\mathcal{M}_{3}\cup\mathcal{M}_{5}$, then the corresponding linear code is not an NMDS code. For the remaining cases, we only need consider the number of nonzero vectors $u\in\mathbb{F}^{4}_{q}$ such that $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=4$. Suppose that a nonzero vector $u\in\mathbb{F}^{4}_{q}$ satisfies this condition. Then $\mathbf{g}_{q+2}\in\mathcal{H}_{u}$ and so $|\mathcal{H}_{u}\cap\mathcal{S}_{h}|=3$. By Lemma 4.25, all planes containing exactly three points of $\mathcal{S}_{h}$ form a $K_{h}$-orbit $\pi(\mathcal{M}_{3})$. This means the number of nonzero vectors $u\in\mathbb{F}^{4}_{q}$ such that $|\mathcal{H}_{u}\cap\mathcal{S}_{G}|=4$ is equal to q1q-1 times the number of planes in $\pi(\mathcal{M}_{3})$ through the point $\mathbf{g}_{q+2}$. By [9, Table 1], these numbers are all great than 0 and only depend on the $K_{h}$-orbit on points to which $\mathbf{g}_{q+2}$ belongs. Thus, the remaining linear codes are all NMDS codes. Obviously, the codes corresponding to $\mathbf{g}_{q+2}$ in the same $K_{h}$-orbit are monomially equivalent, while the codes corresponding to $\mathbf{g}_{q+2}$ in different $K_{h}$-orbits are not monomially equivalent as they have different numbers of minimum weight codes. Therefore, there are two or three equivalence classes of NMDS codes in $\mathscr{C}$ depending on whether $q\equiv 1\pmod{3}$ or $q\equiv-1\pmod{3}$. Finally, by Lemma 2.2, the weight distributions of these codes can be fully determined and we provide them in Table 2. ∎
@@
Let $\mathbf{g}^{\mathrm{T}}_{q+2}=(0,0,1,0)\in\mathcal{M}_{2}$ in (4.1). Then we can directly obtain the corresponding result from [12, Theorem 15] using Theorem 4.26.
@@ 4.2 $q$ odd
[18, Lemma 21.1.3][4, Theorem 2.2] Let $q\equiv 0\pmod{3}$. The group $K_{q}$ has five orbits on the points of $\mathrm{PG}(3,q)$:
@@
Remark 4.31.
In [8], the authors referred to NMDS codes as NMDS-sets. An NMDS-set is said to be complete if it is maximal with respect to set inclusion. They showed that extending $\mathcal{S}_{1}$ by adding a point on a tangent line or an imaginary chord yields an NMDS-set, which can be completed by adding at most one or three additional points, respectively.
@@ 4.3 NMDS codes with parameters [q+3,4,q1][q+3,4,q-1]
where $\mathbb{F}^{*}_{q}=\mathbb{F}_{q}\setminus\{0\}=\{\alpha_{1},\dots,\alpha_{q-1}\}$ and $\mathbf{g}_{q+2},\mathbf{g}_{q+3}\neq\mathbf{0}$. By selecting specific $\mathbf{g}_{q+2}$ and $\mathbf{g}_{q+3}$, we can construct certain NMDS codes. For example, let $p=2$ and $h$ be a positive integer satisfying $\gcd(m,h)=1$, where $m\geq 3$ is odd. Take $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,0,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ or $\mathbf{g}_{q+3}=(0,1,1,0)^{\mathrm{T}}\in\mathcal{M}_{3}$, where $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$ are the point-orbits of $K_{h}$ defined in Lemma 4.24. Then the linear code generated by $G_{3}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code (see [12, Theorems 18 and 19]). A simpler proof of this result can be provided via Theorem 4.26.
@@
Let $h=1$ and $\gcd(3,q-1)=1$. Taking $\mathbf{g}_{q+2}=(0,0,1,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ and $\mathbf{g}_{q+3}=(0,1,0,0)^{\mathrm{T}}\in\mathcal{M}_{2}$ in (4.2), the linear code generated by $G_{3}$ is a [q+3,4,q1][q+3,4,q-1] NMDS code (see [42, Theorem 9]). Similarly to the proof of Theorem 4.32, we further have the following theorem from Tables 2 and 3.
@@
When $q=2^{m}$ with $m$ odd, Theorems 4.32 and 4.34 show that $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$ have identical weight enumerators by Lemma 2.2. However, as noted in [12, Remark 22], $\mathcal{C}_{3}$ and $\mathcal{C}_{5}$ are not monomially equivalent [q+3,4,q1][q+3,4,q-1] NMDS codes for $q=2^{3},2^{5},2^{7}$ and some possible values of $h$. Finally, we have the following proposition.